OpenStudy (anonymous):

Use basic identities to simplify the expression. sin2θ + tan2θ + cos2θ Choices are: sec2θ cos3θ sin θ tan2θ

OpenStudy (anonymous):

This was all I got so far. I'm not sure what to do next! 1/csc+sin/cos+1/sec

OpenStudy (debbieg):

Are you sure you've copied the problem down correctly? Even using a graphing tool, I'm not finding that any of those choices are equivalent to the original expression....?

OpenStudy (anonymous):

Yes! Oh wait sorry, all the numbers are supposed to be exponents. sin^2θ + tan^2θ + cos^2θ sec^2θ cos^3θ sin θ tan^2θ

OpenStudy (debbieg):

ah-hah, that will make a big difference... ;)

OpenStudy (debbieg):

so it's \(\sin^2\theta+\tan^2\theta+\cos^2\theta\), right?

OpenStudy (anonymous):

yes!

OpenStudy (debbieg):

Use your Pythagorean identities... specifically: \(\sin^2\theta+\cos^2\theta=1\) And also \(\tan^2\theta+1=\sec^2\theta\) Now that we have the right problem, it should be pretty trivial. :)

OpenStudy (anonymous):

Is it tan^2θ? :O